# t-READi: Transformer-Powered Robust and Efficient Multimodal Inference for Autonomous Driving

### v2

@@ I Introduction
Autonomous driving, with its worldwide developments [61, 71, 66], promises to achieve greater safety, less harmful emissions, increased lane capacity, and reduced travel time [53]. At the core of autonomous driving, the perception capability of autonomous vehicles (AVs) leverages the data collected from various sensors (e.g., camera, lidar, and radar) to better understand AVs’ surrounding [17], thus enabling applications such as object detection  [77, 28, 20] and semantic segmentation  [18, 11, 59]. To power the perception capability, deep neural networks (DNN) are adopted to process and fuse the multimodal data [40] into a unified representation. Therefore, maintaining the effectiveness of these AV-DNN models is crucial to autonomous driving.
The lifecycle of an AV-DNN model starts with its design, then it gets pre-trained model by individual manufacturers, before being deployed to respective AVs equipped multimodal sensors for performing perception-related inferences. Although DNN-based multimodal fusion for on-device inference in general has attracted attention from both academia [19, 21] and industry (e.g., Google [39], as well as Intel and Ford [24]), only part of them have targeted multimodal DNN for autonomous driving (e.g., [46, 1]). Since this latter batch of proposals largely focuses on DNN architecture design at the manufacturer side, they often disregard the per-vehicle variations in terms of input data and modality, resulting in a missing link, pertaining to in-vehicle inference, between DNN models and their practical adoptions.
@@
One critical issue with existing proposals is the fixed set of multimodal sensors and environment conditions used for model training. In practice, both sensor modalities and environmental conditions may experience drastic variations. For example, different lighting conditions may force the camera to change its exposure, different velocities can result in motion blurs to various degrees, and adverse weather conditions (e.g., rain, fog, and snow) often scatter laser light, thus forcing lidar to perform more intensive scans. Consequently, those two modalities introduce more drastic variations. Moreover, sensor occlusion and/or malfunction may lead to missing modalities that compromise the multimodal fusion. All these realistic scenarios yield variations in the input data and modality, as illustrated in Figure 1. They often significantly degrade the performance of well-designed DNN models, causing erroneous decision-making and even traffic accidents. Therefore, it is crucial to address adverse external conditions and internal sensor malfunctions carefully. As highlighted in the survey by [62], adverse weather conditions and faulty sensors are common contributors to autonomous vehicle accidents, such as the 2019 Tesla Model 3 incident [63]. A naive solution to mitigate performance degradation is to pre-train a large set of models for different scenarios and switch among them during runtime based on the available data and modality variations. However, this solution is not robust against continuously varying sensor and environment conditions since we may only have a countable number of models. Besides, model switching incurs excessive memory footprints and intolerable inference latency.
@@
To address these challenges, we propose t-READi (i.e., transformer-powered Robust and Efficient multimodal inference for Autonomous Driving) as a novel AV inference system; it adaptively accommodates the variations in multimodal sensory data and missing modalities. Specifically, t-READi employs a variation-aware model adaptation algorithm to handle data and modality variations under memory constraints while producing multiple variants of models deviated from the pre-trained model. For each variant, we exploit the parameter-efficient fine-tuning techniques of Large Language Models (LLMs) [30], but extend to widely-used modules in AV-DNN models in addition to transformer modules in LLMs. For example, t-READi injects trainable rank decomposition matrices into residual blocks leading to a significantly decreased number of trainable parameters. With multiple variants of models and the pre-trained model, we can switch different variants according to current input indications, such as weather conditions.
@@ VIII Conclusion
Taking an important step towards full driving automation, we have proposed t-READi in this paper for robust and efficient multimodal inference for autonomous driving. Employing a novel partial weight adaptation mechanism and a data imputation method by autoencoder, t-READi gracefully handles the heterogeneous data caused by sensor parameter variation and missing modality, thus releasing its full potential in the vehicle detection task. With extensive experiments under highly heterogeneous scenarios and comparisons with other baselines, we demonstrate the promising performance of t-READi in vehicle detection for autonomous driving. Overall, we believe that t-READi represents an important step forward in the development of robust and efficient multimodal inference for autonomous driving.

### v3

@@ I Introduction
Autonomous driving, with its worldwide developments [1, 2, 3], promises to achieve greater safety, less harmful emissions, increased lane capacity, and reduced travel time [4, 5, 6]. At the core of autonomous driving, the perception capability of autonomous vehicles (AVs) leverages the data collected from various sensors (e.g., camera, lidar, and radar) to better understand AVs’ surrounding [7, 8, 9], thus enabling applications such as object detection  [10, 11, 12] and semantic segmentation  [13, 14, 15]. To power the perception capability, deep neural networks (DNN) are adopted to process and fuse the multimodal data [16, 17, 18] into a unified representation. Therefore, maintaining the effectiveness of these AV-DNN models is crucial to autonomous driving.
The lifecycle of an AV-DNN model starts with its design, then it gets pre-trained model by individual manufacturers, before being deployed to respective AVs equipped multimodal sensors for performing perception-related inferences. Although DNN-based multimodal fusion for on-device inference in general has attracted attention from both academia [19, 20, 21, 22] and industry (e.g., Google [23], as well as Intel and Ford [24]), only part of them have targeted multimodal DNN for autonomous driving (e.g., [25, 26]). Since this latter batch of proposals largely focuses on DNN architecture design at the manufacturer side, they often disregard the per-vehicle variations in terms of input data and modality, resulting in a missing link, pertaining to in-vehicle inference, between DNN models and their practical adoptions.
@@
One critical issue with existing proposals is the fixed set of multimodal sensors and environment conditions used for model training. In practice, both sensor modalities and environmental conditions may experience drastic variations [27, 28, 29, 30]. For example, different lighting conditions may force the camera to change its exposure, different velocities can result in motion blurs to various degrees, and adverse weather conditions (e.g., rain, fog, and snow) often scatter laser light, thus forcing lidar to perform more intensive scans. Consequently, those two modalities introduce more drastic variations. Moreover, sensor occlusion and/or malfunction may lead to missing modalities that compromise the multimodal fusion. All these realistic scenarios yield variations in the input data and modality, as illustrated in Figure 1. They often significantly degrade the performance of well-designed DNN models, causing erroneous decision-making and even traffic accidents. Therefore, it is crucial to address adverse external conditions and internal sensor malfunctions carefully. As highlighted in the survey by [31], adverse weather conditions and faulty sensors are common contributors to autonomous vehicle accidents, such as the 2019 Tesla Model 3 incident [32]. A naive solution to mitigate performance degradation is to pre-train a large set of models for different scenarios and switch among them during runtime based on the available data and modality variations. However, this solution is not robust against continuously varying sensor and environment conditions since we may only have a countable number of models. Besides, model switching incurs excessive memory footprints and intolerable inference latency.
@@
To address these challenges, we propose t-READi (i.e., transformer-powered Robust and Efficient multimodal inference for Autonomous Driving) as a novel AV inference system; it adaptively accommodates the variations in multimodal sensory data and missing modalities. Specifically, t-READi employs a variation-aware model adaptation algorithm to handle data and modality variations under memory constraints while producing multiple variants of models deviated from the pre-trained model. For each variant, we exploit the parameter-efficient fine-tuning techniques of Large Language Models (LLMs) [34], but extend to widely-used modules in AV-DNN models in addition to transformer modules in LLMs [35, 36, 37]. For example, t-READi injects trainable rank decomposition matrices into residual blocks leading to a significantly decreased number of trainable parameters. With multiple variants of models and the pre-trained model, we can switch different variants according to current input indications, such as weather conditions.
@@ VIII Conclusion
Taking an important step towards full driving automation, we have proposed t-READi in this paper for robust and efficient multimodal inference for autonomous driving. Employing a novel partial weight adaptation mechanism and a data imputation method by autoencoder, t-READi gracefully handles the heterogeneous data caused by sensor parameter variation and missing modality, thus releasing its full potential in the vehicle detection task. With extensive experiments under highly heterogeneous scenarios and comparisons with other baselines, we demonstrate the promising performance of t-READi in vehicle detection for autonomous driving. Overall, we believe that t-READi represents an important step forward in the development of robust and efficient multimodal inference for autonomous driving. As a potential future direction, we are looking forward to extending our t-READ to various applications such as distributed learning systems [96, 97, 98, 99, 100, 101], multi-agent collaborative intelligence [102, 103, 104, 105].